# DDAD: A Two-pronged Adversarial Defense Based on Distributional Discrepancy

### v1

@@ 6 Related Work
We briefly review the related work here, and a more detailed version can be found in Appendix F.
@@ 7 Discussions on Batch-wise Detections
We briefly discuss the limitation, solutions and practicability here, and see Appendix G for detailed discussions.
Limitation and its solutions for user inference. DDAD leverages statistics based on distributional discrepancies, which requires the data to be processed in batches for adversarial detection. When the batch size is too small, the stability of DDAD will be affected (see Figure 2). To address this issue, each single sample provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. We leave finding such statistics as future work.
@@ 8 Conclusion
SADD-based defense methods empirically show that leveraging the distributional discrepancies can effectively defend against adversarial attacks. However, a potential limitation of SADD-based methods is that they will discard data batches that contain AEs, leading to the loss of clean information. To solve this problem, inspired by our theoretical analysis that minimizing distributional discrepancy can help reduce the expected loss on AEs, we propose a two-pronged adversarial defense called Distributional-Discrepancy-based Adversarial Defense (DDAD) that leverages the effectiveness of SADD-based methods without discarding input data. Extensive experiments demonstrate the effectiveness of DDAD against various adversarial attacks. In general, we hope this simple yet effective method could open up a new perspective on adversarial defenses.
@@
This study on adversarial defense mechanisms raises important ethical considerations that we have carefully addressed. We have taken steps to ensure our adversarial defense method is fair. We use widely accepted public benchmark datasets to ensure comparability of our results. Our evaluation encompasses a wide range of attack types and strengths to provide a comprehensive assessment of our defense mechanism. We have also carefully considered the broader impacts of our work. The proposed defense algorithm contributes to the development of more robust machine learning models, potentially improving the reliability of AI systems in various applications. We will actively engage with the research community to promote responsible development and use of adversarial defenses.
@@ Appendix B Mathematical Notations in Section 4
Appendix C Algorithms
3:  for epoch $=1,...,T$ do
6:     $k_{\omega}\leftarrow$ kernel function with parameters $\omega$ using Eq. 4;
7:     $M(\omega)\leftarrow\widehat{\textsc{MMD}}^{2}_{\rm u}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. 3;
8:     $V_{\lambda}(\omega)\leftarrow\hat{\sigma}_{\lambda}(S^{\prime}_{\mathcal{C}},S^{\prime}_{\mathcal{A}};k_{\omega})$ using Eq. 4.2;
9:     $\hat{J}_{\lambda}(\omega)\leftarrow M(\omega)/\sqrt{V_{\lambda}(\omega)}$ using Eq. 4.2;
11:  end for
12:  Output: $k^{*}_{\omega}$
Algorithm 1 Optimizing MMD (Liu et al., 2020).
3:  for epoch $=1,...,T$ do
5:     $S^{\prime}_{\mathcal{A}}\leftarrow$ adversarial examples generated from $(S^{\prime}_{\mathcal{C}},Y^{\prime}_{\mathcal{C}})$;
6:     generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$;
7:     $S^{\prime}_{\text{noise}}\leftarrow S^{\prime}_{\mathcal{A}}+\mathbf{n}$;
[…]
9:     $\bm{\theta}\leftarrow\bm{\theta}-\eta\nabla_{\text{Adam}}(\text{MMD-OPT}(S^{\prime}_{\mathcal{C}},g_{\bm{\theta}}(S^{\prime}_{\text{noise}}))+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S^{\prime}_{\text{noise}})),Y^{\prime}_{\mathcal{C}}))$ using Eq. 4.2;
10:  end for
11:  Output: denoiser $g$ with well-trained parameters $\bm{\theta}^{*}$
Algorithm 2 Training the denoiser.
Algorithm 3 Adaptive white-box PGD+EOT attack for DDAD.
3:  for PGD iteration $1,...,T$ do
[…]
6:     for EOT iteration $1,...,K$ do
[…]
9:        else
10:           Generate Gaussian noise: $\mathbf{n}\sim\mathbb{N}(\mu,\sigma^{2})$; $S_{\text{noise}}\leftarrow S_{\mathcal{A}}+\mathbf{n}$;
11:           ${\mathcal{G}}_{\text{EOT}}\leftarrow{\mathcal{G}}_{\text{EOT}}+\nabla_{S_{\mathcal{A}}}(\text{MMD-OPT}(S_{\mathcal{C}},S_{\mathcal{A}})+\alpha\cdot{\mathcal{L}}_{\text{ce}}(\widehat{h_{\mathcal{C}}^{*}}(g_{\bm{\theta}}(S_{\text{noise}})),Y_{\mathcal{C}}))$;
12:        end if
13:     end for
14:     ${\mathcal{G}}_{\text{EOT}}\leftarrow\frac{1}{K}{\mathcal{G}}_{\text{EOT}}$;
15:     Update adversarial data $S_{\mathcal{A}}\leftarrow\Pi_{\mathcal{B}_{\epsilon}(S_{\mathcal{C}})}\left(S_{\mathcal{A}}+\eta\cdot\text{sign}({\mathcal{G}}_{\text{EOT}})\right)$;
16:  end for
17:  Output: $S_{\mathcal{A}}$
Appendix D Detailed Experiment Settings
D.1 Dataset and target models
@@
We evaluate the effectiveness of DDAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) (small scale) and ImageNet-1K (Deng et al., 2009) (large scale). Specifically, CIFAR-10 contains 50,000 training images and 10,000 test images, divided into 10 classes. ImageNet-1K is a large-scale dataset that contains 1,000 classes and consists of 1,281,167 training images, 50,000 validation images, and 100,000 test images. For the target models, we use three widely used architectures with different scales: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021). Specifically, following Lee & Kim (2023), we use WideResNet-28-10 and WideResNet-70-16 to evaluate the performance of defense methods on CIFAR-10 and we use ResNet-50 to evaluate the performance of defense methods on ImageNet-1K. Additionally, we examine the transferability of our method under different threat models, which include ResNet-18, ResNet-50, WideResNet-70-16 and Swin Transformer.
D.2 Baseline Settings
@@
DDAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DDAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
D.3 Evaluation Settings
@@
We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DDAD. To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. The algorithmic descriptions of the adaptive white-box attack is provided in Algorithm 3. On CIFAR-10, $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks. We also evaluate our method against BPDA+EOT (Hill et al., 2021) on CIFAR-10. For BPDA+EOT, we use the implementation of Hill et al. (2021) with default hyperparameters for evaluation. For transferability experiments, we use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) for evaluation. Specifically, the iteration number of C&W $\ell_{2}$ is set to 200. For $\ell_{\infty}$-norm transfer attacks, we examine the robustness of our method under $\epsilon=8/255$ and $\epsilon=12/255$. For C&W $\ell_{2}$, we examine our method under $\epsilon=0.5$ and $\epsilon=1.0$.
D.4 Implementation Details of DDAD
@@
To avoid the evaluation bias caused by learning similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using the MMA attack with $\ell_{\infty}$-norm (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DDAD. We set $\epsilon=8/255$ with a step size of $2/255$ for CIFAR-10, and $\epsilon=4/255$ with a step size of $1/255$ for ImageNet-1K. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the initial learning rate to $1\times 10^{-3}$ for both CIFAR-10 and ImageNet-1K. We set the epoch number to be 60 and divide the learning rate by 10 at the 45th epoch and 60th epoch to avoid robust overfitting (Rice et al., 2020). The training batch size is set to 500 for CIFAR-10 and 128 for ImageNet-1K. The optimizer we use is Adam (Kingma & Ba, 2015). To improve the training efficiency on ImageNet-1K, we randomly select 100 samples from each class, resulting in 100,000 training samples in total. Notably, during the inference time, we evaluate our method using the entire testing set for both CIFAR-10 and ImageNet-1K. The batch size for evaluation is set to 100 for all datasets.
Appendix E Additional Experiments
E.1 Defending against BPDA+EOT Attack
@@
Table 3: Clean (%) and robust accuracy (%) of defense methods against BPDA+EOT attack under $\ell_{\infty}(\epsilon=8/255)$ threat models on CIFAR-10. We report averaged results and standard deviations of DDAD for 5 runs. We show the most successful defense in bold.
@@
Ours WRN-28-10 DDAD 94.16 $\pm$ 0.08 87.13 $\pm$ 1.19 90.65
E.2 Defending against Unseen Transfer Attacks
Table 4: Robust accuracy (%) of DDAD trained on WideResNet-28-10 against unseen transfer attacks on CIFAR-10. Attackers cannot access parameters of WideResNet-28-10, thereby it is in a gray-box setting. We report averaged results and standard deviations of 5 runs.
Trained on WRN-28-10
Unseen Transfer Attack WRN-70-16 RN-18 RN-50 Swin-T
E.3 Ablation Study on Mixed Data Batches
@@
E.4 Ablation Study on Threshold Values of MMD-OPT
Table 6: Sensitivity of DDAD to the threshold values of MMD-OPT on CIFAR-10. We report clean and robust accuracy (%) against adaptive white-box attacks ($\epsilon=8/255$). The classifier used is WRN-28-10.
@@ E.4 Ablation Study on Threshold Values of MMD-OPT
Table 7: Sensitivity of DDAD to the threshold values of MMD-OPT on ImageNet-1K. We report clean and robust accuracy (%) against adaptive white-box attacks ($\epsilon=4/255$). The classifier used is RN-50.
Threshold Value Clean $\ell_{\infty}$)
0.010 76.61 53.75
0.015 76.61 53.75
0.020 78.61 53.75
0.025 78.61 53.75
0.030 78.61 0.46
0.040 78.61 0.46
0.050 78.61 0.46
E.5 Ablation Study on Injecting Gaussian Noise
@@
E.6 Ablation Study on the Two-pronged Process
@@
E.7 Combination with Adversarial Training

### v2

@@ 6 Related Work
We briefly review the related work here, and a more detailed version can be found in Appendix E.
@@ 7 Discussions on Batch-wise Detections
We briefly discuss the limitation, solutions and practicability here, and see Appendix F for detailed discussions.
Limitation and its solutions for user inference. DAD leverages statistics based on distributional discrepancies, which requires the data to be processed in batches for adversarial detection. When the batch size is too small, the stability of DAD will be affected (see Figure 2). To address this issue, each single sample provided by the user can be dynamically stored in a queue. Once the queue accumulates enough samples to form a batch, our method can then process the batch collectively using the proposed approach. A direct cost of this solution is the waiting time, as the system must accumulate enough samples (e.g., 50 samples) to form a batch before processing. However, in scenarios where data arrives quickly, the waiting time is typically very short, making this approach feasible for many real-time applications. Overall, it is a trade-off problem: using our method for user inference can obtain high robustness, but the cost is to wait for batch processing. Based on the performance improvements our method obtains over the baseline methods, we believe the cost is acceptable. Another possible solution is to find more robust statistics that can measure distributional discrepancies with fewer samples. We leave finding such statistics as future work.
@@ 8 Conclusion
SADD-based defense methods empirically show that leveraging the distributional discrepancies can effectively defend against adversarial attacks. However, a potential limitation of SADD-based methods is that they will discard data batches that contain AEs, leading to the loss of clean information. To solve this problem, inspired by our theoretical analysis that minimizing distributional discrepancy can help reduce the expected loss on AEs, we propose a two-pronged adversarial defense called Distributional-discrepancy-based Adversarial Defense (DAD) that leverages the effectiveness of SADD-based methods without discarding input data, which kills two birds with one stone. Extensive experiments demonstrate that DAD effectively defends against various adversarial attacks, simultaneously improving both robustness and clean accuracy. In general, we hope this simple yet effective method could open up a new perspective on adversarial defenses based on distributional discrepancies.
Acknowledgements
JCZ is supported by the Melbourne Research Scholarship and would like to thank Qiwei Tian and Ruijiang Dong for productive discussions. FL is supported by the Australian Research Council (ARC) with grant number DE240101089, LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037. BR is supported by ARC DP220102269. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ Acknowledgements
This study on adversarial defense mechanisms raises important ethical considerations that we have carefully addressed. We have taken steps to ensure our adversarial defense method is fair. We use widely accepted public benchmark datasets to ensure comparability of our results. Our evaluation encompasses a wide range of attack types and strengths to provide a comprehensive assessment of our defense mechanism. The proposed defense algorithm contributes to the development of more robust machine learning models, potentially improving the reliability of AI systems in various applications. We will actively engage with the research community to promote responsible development and use of adversarial defenses.
@@ Appendix B Mathematical Notations in Section 4
Appendix C Detailed Experiment Settings
C.1 Dataset and target models
@@
We evaluate the effectiveness of DAD on two benchmark datasets with different scales, i.e., CIFAR-10 (Krizhevsky et al., 2009) (small scale) and ImageNet-1K (Deng et al., 2009) (large scale). Specifically, CIFAR-10 contains 50,000 training images and 10,000 test images, divided into 10 classes. ImageNet-1K is a large-scale dataset that contains 1,000 classes and consists of 1,281,167 training images, 50,000 validation images, and 100,000 test images. For the target models, we use three widely used architectures with different scales: ResNet (He et al., 2016), WideResNet (Zagoruyko & Komodakis, 2016) and Swin Transformer (Liu et al., 2021b). Specifically, following Lee & Kim (2023), we use WideResNet-28-10 and WideResNet-70-16 to evaluate the performance of defense methods on CIFAR-10 and we use ResNet-50 to evaluate the performance of defense methods on ImageNet-1K. Additionally, we examine the transferability of our method under different threat models, which include ResNet-18, ResNet-50, WideResNet-70-16 and Swin Transformer.
C.2 Baseline Settings
@@
DAD is a two-pronged adversarial defense method, which is different from most existing defense methods. In terms of the pipeline structure, MagNet (Meng & Chen, 2017) is the only similar defense method to ours, which also contains a two-pronged process. However, MagNet is now considered outdated, making it unfair for DAD to compare with it. Therefore, to make the comparison as fair as possible, we follow a recent study on robust evaluation (Lee & Kim, 2023) to compare our method with SOTA adversarial training (AT) methods in RobustBench (Croce et al., 2020) and adversarial purification (AP) methods selected by Lee & Kim (2023).
C.3 Evaluation Settings
@@
We mainly use PGD+EOT (Athalye et al., 2018b) and AutoAttack (Croce & Hein, 2020a) to compare our method with different baseline methods. Specifically, following Lee & Kim (2023), we evaluate AP methods on the PGD+EOT attack with 200 PGD iterations for CIFAR-10 and 20 PGD iterations for ImageNet-1K. We set the EOT iteration to 20 for both datasets. We evaluate AT baseline methods using AutoAttack with 100 update iterations, as AT methods have seen PGD attacks during training, leading to overestimated results when evaluated on PGD+EOT (Lee & Kim, 2023). For our method, we implicitly design an adaptive white-box attack by considering the entire defense mechanism of DAD. To make a fair comparison, we evaluate our method on both adaptive white-box PGD+EOT attack and adaptive white-box AutoAttack with the same configurations mentioned above. Notably, we find that our method achieves the worst-case robust accuracy on adaptive white-box PGD+EOT attack. Therefore, we report the robust accuracy of our method on adaptive white-box PGD+EOT attack for Table 1 and 2. The algorithmic descriptions of the adaptive white-box attack is provided in Algorithm 3. On CIFAR-10, $\epsilon$ for $\ell_{\infty}$-norm-based attacks and $\ell_{2}$-norm-based attacks is set to $8/255$ and $0.5$, respectively. While on ImageNet-1K, we set $\epsilon=4/255$ for $\ell_{\infty}$-norm-based attacks. We also evaluate our method against BPDA+EOT (Hill et al., 2021) on CIFAR-10. For BPDA+EOT, we use the implementation of Hill et al. (2021) with default hyperparameters for evaluation. For transferability experiments, we use PGD+EOT $\ell_{\infty}$ and C&W $\ell_{2}$ (Carlini & Wagner, 2017) for evaluation. Specifically, the iteration number of C&W $\ell_{2}$ is set to 200. For $\ell_{\infty}$-norm transfer attacks, we examine the robustness of our method under $\epsilon=8/255$ and $\epsilon=12/255$. For C&W $\ell_{2}$, we examine our method under $\epsilon=0.5$ and $\epsilon=1.0$.
C.4 Implementation Details of DAD
@@
To avoid the evaluation bias caused by learning similar attacks beforehand during training, we train both the MMD-OPT and the denoiser using the MMA attack with $\ell_{\infty}$-norm (Gao et al., 2022), which differs significantly from PGD+EOT and AutoAttack. Then, we use unseen attacks to evaluate DAD. We set $\epsilon=8/255$ with a step size of $2/255$ for CIFAR-10, and $\epsilon=4/255$ with a step size of $1/255$ for ImageNet-1K. For optimizing the MMD, following Gao et al. (2021), we set the learning rate to be $2\times 10^{-4}$ and the epoch number to be 200. For training the denoiser, we set the initial learning rate to $1\times 10^{-3}$ for both CIFAR-10 and ImageNet-1K. We set the epoch number to be 60 and divide the learning rate by 10 at the 45th epoch and 60th epoch to avoid robust overfitting (Rice et al., 2020). The training batch size is set to 500 for CIFAR-10 and 128 for ImageNet-1K. The optimizer we use is Adam (Kingma & Ba, 2015). To improve the training efficiency on ImageNet-1K, we randomly select 100 samples from each class, resulting in 100,000 training samples in total. Notably, during the inference time, we evaluate our method using the entire testing set for both CIFAR-10 and ImageNet-1K. The batch size for evaluation is set to 100 for all datasets.
Appendix D Additional Experiments
D.1 Defending against BPDA+EOT Attack
@@
Table 4: Clean (%) and robust accuracy (%) of defense methods against BPDA+EOT attack under $\ell_{\infty}(\epsilon=8/255)$ threat models on CIFAR-10. We report averaged results and standard deviations of DAD for 5 runs. We show the most successful defense in bold.
@@
Ours WRN-28-10 DAD 94.16 $\pm$ 0.08 87.13 $\pm$ 1.19 90.65
D.2 Ablation Study on Mixed Data Batches
@@
D.3 Ablation Study on Threshold Values of MMD-OPT
Table 6: Sensitivity of DAD to the threshold values of MMD-OPT on CIFAR-10. We report clean and robust accuracy (%) against adaptive white-box attacks ($\epsilon=8/255$). The classifier used is WRN-28-10.
@@ D.3 Ablation Study on Threshold Values of MMD-OPT
D.4 Ablation Study on the Regularization Term $\alpha$ in Eq. (4.2)
Table 7: Sensitivity of DAD to the regularization term $\alpha$ on CIFAR-10. We report clean and robust accuracy (%) against adaptive white-box PGD+EOT ($\epsilon=8/255$). The classifier used is WRN-28-10.
$\alpha$ Clean PGD+EOT
0.01 94.16 67.53
0.05 94.16 50.70
0.10 94.16 45.35
D.5 Ablation Study on Injecting Gaussian Noise
@@
D.6 Ablation Study on the Two-pronged Process
@@
D.7 Compute Resources
Table 10: Training time (hours : minutes : seconds) and memory consumption (MB) for DAD on CIFAR-10 and ImageNet-1K . This table reports the compute resources for the entire training process of DAD (i.e., optimizing MMD + training the denoiser).
Dataset GPU Batch Size Classifier Training Time Memory
CIFAR-10 2 $\times$ NVIDIA A100 500 RN-18 00:28:17 5927
WRN-28-10 00:55:34 6276
ImageNet-1K 4 $\times$ NVIDIA A100 128 RN-50 09:52:50 97354
Table 11: Inference time (hours : minutes : seconds) for DAD on CIFAR-10 and ImageNet-1K. This table reports the comput resources for evaluating the entire test set of CIFAR-10 (i.e., 10,000 images) and ImageNet-1K (i.e., 50,000 images).
Dataset GPU Batch Size Classifier Inference Time
CIFAR-10 1 $\times$ NVIDIA A100 100 WRN-28-10 00:00:32
ImageNet-1K 2 $\times$ NVIDIA A100 100 RN-50 00:03:08
Table 10 presents the compute resources for DAD, which include GPU configurations, batch size, classifier, training time, and memory usage for each dataset. For CIFAR-10, using 2 NVIDIA A100 GPUs with a batch size of 500, our method’s training time is approximately 28 minutes with ResNet-18 and 55 minutes with WideResNet-28-10. The memory consumption is 5927 MB and 6276 MB, respectively. For ImageNet-1K, using 4 NVIDIA A100 GPUs with a batch size of 128, our method’s training time is approximately 10 hours, with a memory consumption of 97354 MB. Compared to AT baseline methods, DAD offers better training efficiency (e.g., it can scale to large datasets like ImageNet-1K). This is mainly because we directly use the pre-trained classifier. Furthermore, training MMD is extremely fast (usually less than 1 minute on CIFAR-10) and we use a lightweight denoiser.
Table 11 presents the compute resources for evaluating DAD, which include GPU configurations, batch size, classifier and inference time for each dataset. For CIFAR-10, using 1 NVIDIA A100 GPU with a batch size of 100, our method’s inference time is approximately 32 seconds over the entire test set of CIFAR-10. For ImageNet-1K, using 2 NVIDIA A100 GPUs with a batch size of 100, our method’s inference time is approximately 3 minutes over the entire test set of ImageNet-1K. Although DAD requires training an extra denoiser and MMD-OPT, it significantly outperforms AP baselines in inference speed. Furthermore, relying on a pre-trained generative model is not always feasible, as training such models at scale can be highly resource-intensive. Therefore, considering considering the trade-off between computational cost and the performance of DAD, we believe that training an additional detector and denoiser is feasible and worthwhile.
D.8 Combination with Adversarial Training